# Blink: Dynamic Visual Token Resolution for Enhanced Multimodal Understanding

### v1

@@ 1 Introduction
Multimodal large language models (MLLMs) have rapidly emerged across various downstream fields, including computer vision and natural language processing [1, 41, 4, 10, 26]. By coupling a vision encoder with a projector that maps visual features into the language space of the following LLM, MLLMs effectively unify visual and textual representations [2, 28, 32]. This architecture has demonstrated impressive performance across a wide spectrum of multimodal tasks, including text generation, visual understanding, and cross-modal reasoning. Despite these advances, MLLMs still struggle with insufficient visual perception capabilities [45, 11], often resulting in incorrect answers or hallucinated explanations [42, 5, 22, 15]. Compared with LLMs, MLLMs take additional visual inputs and thus need to reason over both linguistic and visual information, making their performance on multimodal tasks highly dependent on visual perception. However, current MLLMs still adopt the conventional LLM architecture for downstream task execution, lack of explicitly enhancing or leveraging salient visual regions [46, 6, 52]. Therefore, it is imperative to investigate their visual perception to reveal potential risks in downstream applications.
@@ 2.1 Architecture of MLLMs
Figure 2: Visualization of attention maps across layers. In the right panels, lighter colors indicate higher attention from the last text token, and red boxes highlight the region with the highest total attention weights.
Figure 3: Attention ratio across layers. Purple lines indicate the ratio of attention weights in the most attended region to the total attention over all visual tokens, and red dots mark layers where the most attended region is correct.
@@ 3.2.2 Token Drop
Table 1: Downstream task performance across multiple benchmarks on LLaVA-1.5. Results of an additional MLLM backbone are provided in Appendix C. Vanilla denotes the base model, and our methods correspond to two configurations of Blink, where -interp indicates the variant that replaces the amplifier with training-free interpolation while retaining the Blink inference pipeline. The best scores are in bold.
@@ 4.3 Ablation Studies
Key Modules. To validate the effectiveness of our proposed modules, we perform ablation studies using MME benchmark on the two key components of Blink: saliency-guided scanning (SGS) and dynamic token resolution (DTR). In the w/o SGS variant, token selection for expansion or drop is performed randomly, without guidance of saliency maps. In the w/o DTR variant, token operations follow a fixed expand–drop cycle across layers, without dynamically adjusting based on thresholds. As shown in Table 2, removing either SGS or DTR results in notable performance drops for both Blink and its variant Blink-interp. Specifically, removing SGS results in the total score dropping by 2.10 for Blink-interp and 2.15 for Blink, and removing DTR causes a larger decrease of 8.63 for Blink-interp and 41.07 for Blink, underscoring the critical role of both saliency-guided scanning and dynamic token resolution in effective token selection and capturing salient visual information.
@@
Ratio Thresholds. We further investigate the impact of different expansion and drop thresholds on performance. Three alternative settings to our main configuration are considered: (1) the extreme case of disabling token drop (w/o Drop), (2) high expansion threshold and low drop threshold (High $\tau_{\text{exp}}$ / Low $\tau_{\text{drop}}$), and (3) high expansion threshold only (High $\tau_{\text{exp}}$), with the drop threshold kept at its default value. As shown in Table 2, these settings lead to varying performance compared to our default configuration. Specifically, for Blink-interp, the High $\tau_{\text{exp}}$ setting yields only a marginal improvement of 0.40 in the total score, whereas the w/o Drop and High $\tau_{\text{exp}}$ / Low $\tau_{\text{drop}}$ settings reduce performance by 1.39 and 1.75, respectively. For Blink, disabling token drops increases performance by 2.50, while the High $\tau_{\text{exp}}$ / Low $\tau_{\text{drop}}$ and High $\tau_{\text{exp}}$ settings decrease it by 13.84 and 15.99. These results indicate that completely disabling token deletion leads to divergent outcomes, while excessively high expansion thresholds limit effective token expansion, demonstrating that both token expansion and token drop contribute to our framework.
Figure 6: Visualization of attention redistribution after token expansion. Red boxes in the original image indicate the ground-truth important regions. The right panels show attention distributions on the original and expanded visual tokens at layers 13 and 20. See Appendix D for more visualization results.
4.4 Analysis of Attention Redistribution
In this section, we investigate how tokens generated by the token super-resolution module influence attention redistribution across transformer layers after expansion. Understanding this process is important to verify whether the expanded tokens effectively guide the model to focus and contribute to improved visual perception. For clarity, expansion is applied at layer 12, and the new tokens are retained in subsequent layers without dropping.
Fig. 6 shows attention maps at different layers to illustrate this spatial redistribution. In both cases, attention on the new tokens is more evenly distributed than on the original tokens, with their total attention consistently higher, indicating effective focus on the expanded tokens. In Case 2, attention initially targets an incorrect region at layer 13 but shifts to the correct region by layer 20, which is consistent with the insight in Sec. 2.2 that the model can progressively adjust attention across layers.
4.5 Analysis of Layer Range
We observe an interesting phenomenon that although selecting intermediate layers (e.g., our 12–18 layer range) generally yields good performance, slight shifts in the chosen layer range can lead to measurable differences in results. As shown in Table 3, we compare several layer selection strategies on the two types of tasks in the MME benchmark as well as the total performance. Shifting from 12–18 to 11–17 slightly improves the total score by 0.96, while shifting to 13–19 or 14–20 results in moderate decreases of 7.42 and 7.33, respectively. These results indicate that different layer ranges contribute differently to the overall performance, and even small changes can lead to non-negligible variations.
To better understand this effect, we further evaluate the importance and contribution of each layer through three analyses: (1) varying the starting layer while fixing the ending layer at 18, (2) fixing the starting layer at 12 while varying the ending layer, and (3) performing dynamic token operations on single layers based on thresholds. As shown in Fig. 7, selecting a single layer leads to clear performance differences, suggesting that attention distributions and accuracy vary across layers. In particular, layers around 7–26 produce much stronger improvements, whereas using only early layers around 4–6 leads to substantial performance drops. Moreover, the choice of layer range affects performance. In the first analysis, excluding layers 4–9 yields better results, whereas in the second analysis, including layers 14–18 causes a noticeable drop. These observations are consistent with Sec. 2.2, where correct attention tends to emerge in the middle layers. And, from the curves marked with our config, it can be seen that we adopt a broadly effective layer range, though more fine-grained layer selection may further improve performance.  [^1] More analysis experiments, future directions, and limitations are provided in the supplementary materials.
@@ 4.5 Analysis of Layer Range
Figure 7: Performance changes on the total MME score with different layer ranges. The top panel varies the starting layer up to layer 18, while the bottom panel varies the ending layer from layer 12. Purple lines indicate performance across layer ranges, and yellow lines show performance when dynamic operations are applied to a single layer.
@@ 5 Related Work
Multimodal Large Language Models. Driven by the success of large language models (LLMs), increasing attention has been devoted to developing end-to-end multimodal large language models (MLLMs). Early works such as CLIP [38] and ALIGN [21] aligned visual and textual representations through contrastive learning, laying the foundation for cross-modal understanding. Following this paradigm, MLLMs typically employ a vision encoder and a projector to map visual embeddings into the language space of an LLM. Representative models such as BLIP-2 [28], LLaVA [31], InternVL [10] and the Qwen-VL [3] series adopt this modular design, where projected visual tokens and textual embeddings are jointly processed by the LLM for multimodal tasks. While this unified token-based framework enables generalization across modalities, it overlooks the inherent differences between visual and textual information. As a result, MLLMs often exhibit insufficient visual perception and hallucinated reasoning [42, 5, 22, 15], and their performance on multimodal reasoning tasks remains highly dependent on accurate visual grounding [45, 11]. This motivates further investigation into enhancing the visual perception and grounding capabilities of MLLMs to ensure reliable multimodal understanding.
@@ A.2 Datasets
To train the TokenSR modules, we use the processed LLaVA-1.5 [32] training set, which is constructed from several widely used vision–language datasets. For each image, we additionally generate four quadrant crops (top-left, top-right, bottom-left, bottom-right) and record their positions relative to the original image. Each processed sample therefore consists of the full image, one cropped quadrant, and its corresponding positional metadata.
During training, both the full image and the cropped quadrant are converted into pixel-value tensors and fed through the frozen MLLM backbone to obtain their hidden representations. The cropped image provides the teacher features, while the TokenSR module operates on the full-image tokens to generate expanded tokens. The module is optimized so that these expanded tokens match the teacher features at the corresponding spatial region.
@@ A.3 Hyper-parameters
The training hyper-parameters for the TokenSR modules are presented in Table 4. All experiments are conducted on 8 H800 (80G) GPUs in total, where each TokenSR module is trained on 2 GPUs with a global batch size of 8, using BFloat16 precision. We set the warmup ratio to 3% and employ a cosine learning rate scheduler, decaying the learning rate from 1e-4 to 0. Our implementation is based on HuggingFace Transformers [44] and DeepSpeed [39].
@@ Appendix B Implementation Details
Appendix C Evaluation on Other Backbone
We further evaluate Blink on the LLaVA-NeXT-7B [33] backbone. For POPE, MMBench, MMBench-CN, GQA, ScienceQA, and MM-Vet, we use layers 14–20 with $\tau_{\text{exp}}=0.5$ and $\tau_{\text{drop}}=0.4$, while for MME, we use layers 12–20 with thresholds $\tau_{\text{exp}}=0.4$ and $\tau_{\text{drop}}=0.3$. As shown in Table 5, our method consistently improves performance across a wide range of downstream benchmarks. Blink yields a notable gain of 15.00 on MMECognition compared with the vanilla model and achieves the highest scores on GQA, MMBench-CN, POPE, and MM-Vet. Although its MMEPerception score is slightly lower than that of Blink-interp, Blink shows stronger advantages on most tasks, suggesting that the fully trained TokenSR module strengthens fine-grained multimodal understanding.
Blink-interp achieves the best performance on MMEPerception and improves over the backbone model on ScienceQA and MM-Vet, highlighting the complementary benefits of the dynamic inference pipeline. These results collectively demonstrate that Blink generalizes well to stronger backbones and provides robust gains across diverse downstream tasks.
Appendix D Cases of Attention Redistribution
@@
Appendix E Ablation Study on Patch Numbers
As shown in Table 6, neither the $3\times 3$ nor $4\times 4$ partitioning yields improvements over the default $2\times 2$ setting. For Blink-interp, increasing the number of patches consistently degrades both perception and cognition scores, with the $4\times 4$ configuration showing the largest decline. For Blink, the $3\times 3$ variant offers a small gain in cognition but still reduces perception and overall MME performance, while the $4\times 4$ setup leads to drops across all metrics. This may be because the current thresholds are derived using a simple proportional scaling rule, and more fine-grained tuning could further optimize performance for different patch granularities. Nevertheless, across all configurations, Blink equipped with the trained TokenSR module consistently outperforms Blink-interp, confirming that the learned amplifier is more effective than naive interpolation under all patch partitions.
@@ Appendix E Ablation Study on Patch Numbers
Appendix F Ablation Study on Interpolation
Table 7 presents the performance on the MME benchmark with and without the interpolation step. In our experiments, the w/o interp condition refers to bypassing this initial upsampling. Specifically, after selecting the saliency tokens, they are directly extracted and inserted between the original visual and text tokens without resizing. For Blink-interp, w/o interp corresponds to inserting the saliency tokens directly into the dynamic inference pipeline without any training. For Blink, w/o interp means feeding the tokens into the trained TokenSR module without prior interpolation. From the experimental results, skipping interpolation for Blink-interp slightly reduces the perception score by 3.50, indicating that aligning the length of the expanded saliency sequence with the original image helps preserve spatial information. For Blink, removing interpolation leads to drops of 4.66 in perception, 6.79 in cognition, and 11.45 overall, further demonstrating that the TokenSR module benefits from receiving saliency tokens of the same length as the original image tokens, which allows the amplifier to more effectively refine multimodal representations. Overall, these results highlight that interpolation is a crucial step for fully leveraging spatial context and ensuring stable and accurate performance.
@@ Appendix F Ablation Study on Interpolation
Appendix G Limitations and Future Work

### v2

@@ 1 Introduction
Multimodal large language models (MLLMs) have rapidly emerged across various downstream fields, including computer vision and natural language processing [1, 41, 4, 10, 26]. By coupling a vision encoder with a projector that maps visual features into the language space of the following large language models (LLMs), MLLMs effectively unify visual and textual representations [2, 28, 32]. This architecture has demonstrated impressive performance across a wide spectrum of multimodal tasks, including text generation, visual understanding, and cross-modal reasoning. Despite these advances, MLLMs still struggle with insufficient visual perception capabilities [45, 11], often resulting in incorrect answers or hallucinated explanations [42, 5, 22, 15]. Compared with LLMs, MLLMs take additional visual inputs and thus need to reason over both linguistic and visual information, making their performance on multimodal tasks highly dependent on visual perception. However, current MLLMs still adopt the conventional LLM architecture for downstream task execution, lack of explicitly enhancing or leveraging salient visual regions [46, 6, 52]. Therefore, it is imperative to investigate their visual perception to reveal potential risks in downstream applications.
@@ 2.1 Architecture of MLLMs
Figure 2: Visualization of attention maps across layers. In the right panels, lighter colors indicate higher attention from the last text token, and red boxes mark the highest-attention regions.
Figure 3: Attention ratio across layers. Purple lines show the ratio of the most attended region to total visual-token attention, and red dots mark layers where the most attended region is correct.
@@ 3.2.2 Token Drop
Table 1: Downstream task performance across multiple benchmarks on LLaVA-1.5. Results of other backbones are provided in Appendix C. Vanilla denotes the base model, and our methods correspond to two configurations of Blink, where -interp indicates the variant that replaces the amplifier with training-free interpolation while retaining the Blink inference pipeline. The best scores are in bold.
@@ 4.2 Main Results
Figure 6: Visualization of attention redistribution after token expansion. Red boxes in the original image indicate the ground-truth important regions. The right panels show attention distributions on the original and expanded visual tokens at layers 13 and 20. See Appendix E for more visualization results.
@@ 4.3 Ablation Studies
Key Modules. To validate the effectiveness of our proposed modules, we perform ablation studies using MME benchmark on the two key components of Blink: saliency-guided scanning (SGS) and dynamic token resolution (DTR). In the w/o SGS variant, token selection for expansion or drop is performed randomly, without guidance of saliency maps. In the w/o DTR variant, token operations follow a fixed expand–drop cycle across layers, without dynamically adjusting based on thresholds. As shown in Tab. 2, removing either SGS or DTR results in notable performance drops for both Blink and its variant Blink-interp. Specifically, removing SGS results in the total score dropping by 2.10 for Blink-interp and 2.15 for Blink, and removing DTR causes a larger decrease of 8.63 for Blink-interp and 41.07 for Blink, underscoring the critical role of both saliency-guided scanning and dynamic token resolution in effective token selection and capturing salient visual information.
@@
Ratio Thresholds. We further investigate the impact of different expansion and drop thresholds on performance. Three alternative settings to our main configuration are considered: (1) the extreme case of disabling token drop (w/o Drop), (2) high expansion threshold and low drop threshold (High $\tau_{\text{exp}}$ / Low $\tau_{\text{drop}}$), and (3) high expansion threshold only (High $\tau_{\text{exp}}$), with the drop threshold kept at its default value. As shown in Tab. 2, these settings lead to varying performance compared to our default configuration. Specifically, for Blink-interp, the High $\tau_{\text{exp}}$ setting yields only a marginal improvement of 0.40 in the total score, whereas the w/o Drop and High $\tau_{\text{exp}}$ / Low $\tau_{\text{drop}}$ settings reduce performance by 1.39 and 1.75, respectively. For Blink, disabling token drops increases performance by 2.50, while the High $\tau_{\text{exp}}$ / Low $\tau_{\text{drop}}$ and High $\tau_{\text{exp}}$ settings decrease it by 13.84 and 15.99. These results indicate that completely disabling token deletion leads to divergent outcomes, while excessively high expansion thresholds limit effective token expansion, demonstrating that both token expansion and token drop contribute to our framework.
@@
4.4 Analysis of Attention Redistribution
In this section, we investigate how tokens generated by the token super-resolution module influence attention redistribution across transformer layers after expansion. This helps verify whether the expanded tokens guide the model to focus better and improve visual perception. For clarity, expansion is applied at layer 12, and the new tokens are retained in all subsequent layers.
Fig. 6 shows attention maps at different layers to illustrate this spatial redistribution. In both cases, attention over the new tokens is more evenly distributed and consistently higher than that of the original tokens, indicating effective focus. In Case 2, attention initially concentrates on an incorrect region at layer 13 but shifts to the correct region by layer 20, consistent with Sec. 2.2, where attention is progressively refined across layers.
4.5 Analysis of Layer Range
We observe that although selecting intermediate layers (e.g., 12–18) generally yields strong performance, small shifts in the layer range can lead to noticeable differences. As shown in Tab. 3, shifting from 12–18 to 11–17 slightly improves the total score by 0.96, while shifting to 13–19 or 14–20 decreases performance by 7.42 and 7.33, respectively. This suggests that different layer ranges contribute unevenly, and even minor changes can have non-negligible impact.
To better understand this effect, we further evaluate the importance and contribution of each layer through three analyses: (1) varying the starting layer with the ending layer fixed at 18, (2) varying the ending layer with the starting layer fixed at 12, and (3) applying dynamic token operations to individual layers. As shown in Fig. 7, single-layer selection leads to clear performance differences, indicating that attention and accuracy vary across layers. In particular, layers around 7–26 produce much stronger improvements, whereas using only early layers around 4–6 leads to substantial performance drops. Moreover, the choice of layer range affects performance. In the first analysis, excluding layers 4–9 yields better results, whereas in the second analysis, including layers 14–18 causes a noticeable drop. These observations are consistent with Sec. 2.2, where correct attention tends to emerge in the middle layers. And, from the curves marked with our config, it can be seen that we adopt a broadly effective layer range, though more fine-grained layer selection may further improve performance.
Figure 7: Performance changes on the total MME score with different layer ranges. The top panel varies the start layer (up to 18), and the bottom panel varies the end layer (from 12). Purple lines show performance across ranges, and yellow lines show single-layer dynamic performance.
@@ 5 Related Work
Multimodal Large Language Models. Driven by the success of large language models (LLMs), increasing attention has been devoted to developing end-to-end multimodal large language models (MLLMs). Early works such as CLIP [38] and ALIGN [21] aligned visual and textual representations through contrastive learning, laying the foundation for cross-modal understanding. Following this paradigm, MLLMs typically employ a vision encoder and a projector to map visual embeddings into the language space of an LLM. Representative models such as BLIP-2 [28], LLaVA [31], InternVL [10] and the Qwen-VL [3] series adopt this modular design, where projected visual tokens and textual embeddings are jointly processed by the LLM for multimodal tasks. However, this unified token-based paradigm overlooks modality differences, leading to limited visual perception and hallucination issues [42, 5, 22, 15]. Consequently, performance on multimodal reasoning heavily depends on accurate visual grounding [45, 11], motivating further improvements in visual perception.
@@ 6 Conclusion
Acknowledgement
We would like to thank the anonymous reviewers for their comments, which greatly improve our paper. This work is supported by the National Natural Science Foundation of China (No. 62472419, 62472420) and the Enterprise Project (No. E4V06811F3).
@@ A.2 Datasets
To train the TokenSR modules, we use the processed LLaVA-1.5 [32] training set, which is constructed from several widely used vision–language datasets. For each image, we additionally generate four quadrant crops (top-left, top-right, bottom-left, bottom-right) and record their positions relative to the original image. Each processed sample therefore consists of the full image, one cropped image, and the corresponding positional metadata of the cropped quadrant.
During training, both the full image and the cropped image are converted into pixel-value tensors and fed into the frozen MLLM backbone to obtain hidden representations. The cropped image provides the teacher features, while the TokenSR module takes as input the token hidden states corresponding to the same spatial region in the full image after bilinear interpolation, and generates enhanced token hidden states. The optimization objective is to align the enhanced token hidden states with the teacher features of the corresponding spatial region.
@@ A.3 Hyper-parameters
The training hyper-parameters for the TokenSR modules are presented in Tab. 4. All experiments are conducted on 8 H800 (80G) GPUs in total, where each TokenSR module is trained on 2 GPUs with a global batch size of 8, using BFloat16 precision. We set the warmup ratio to 3% and employ a cosine learning rate scheduler, decaying the learning rate from 1e-4 to 0. Our implementation is based on HuggingFace Transformers [44] and DeepSpeed [39].
@@ Appendix B Implementation Details
Appendix C Evaluation on Other Backbones
C.1 Evaluation on LLaVA-NeXT
[…]
C.2 Evaluation on Qwen2.5-VL
[…]
Vanilla 1638.16 598.57 58.14 87.63 80.81
Blink-interp (Ours) 1638.16 606.07 58.37 87.82 81.06
Blink (Ours) 1645.65 608.57 58.28 87.75 81.21
Table 6: Downstream task performance on Qwen2.5-VL-7B. Vanilla denotes the base model, and our methods correspond to two configurations of Blink, where -interp indicates the variant that replaces the amplifier with training-free interpolation while retaining the Blink inference pipeline. The best scores are in bold, and the second-best results are underlined.
To further validate the effectiveness of Blink on modern multimodal backbones, we conduct additional experiments on Qwen2.5-VL-7B [4], a recent and strong vision-language model. We evaluate our method across four benchmarks, including MME [17], GQA [20], POPE [29], and ScienceQA [35]. We consider both Blink and Blink-interp, where the trainable modules are replaced with bilinear interpolation while retaining the same inference pipeline.
[…]
ViCrop 1804.65 60.98 87.25 21.5T 3
Blink (Ours) 1881.53 61.98 85.23 9.73T 1
Table 7: Downstream task performance and efficiency of different methods on LLaVA-1.5-7B. FLOPs denote the theoretical maximum computation per image. #Fwd Passes indicates the number of full model inferences required for each input. The best scores are in bold.
Appendix D Comparison with ViCrop
Direct comparison with dynamic perception methods is non-trivial, as they typically require multiple forward passes and incur higher inference costs. To ensure a fair evaluation, we reproduce ViCrop [51] on LLaVA-1.5-7B and report results on three benchmarks. As shown in Tab. 7, our method achieves better performance on MME and GQA, while maintaining comparable results on POPE. These results demonstrate that Blink consistently improves multimodal perception over prior methods.
In terms of efficiency, Blink significantly reduces computational overhead by lowering FLOPs from 21.5T to 9.73T (a 54.7% reduction) and requiring only a single forward pass instead of three, thereby reducing inference latency. Meanwhile, the additional memory overhead compared to vanilla LLaVA-1.5-7B is minimal, with GPU usage increasing from 19.7 GB to 20.8 GB (a 1.1 GB increase) under FP16. Overall, Blink achieves a better balance between performance and efficiency, making it more suitable for practical deployment.
Appendix E Cases of Attention Redistribution
@@
Appendix F Ablation Study on Patch Numbers
As shown in Tab. 8, neither the $3\times 3$ nor $4\times 4$ partitioning yields improvements over the default $2\times 2$ setting. For Blink-interp, increasing the number of patches consistently degrades both perception and cognition scores, with the $4\times 4$ configuration showing the largest decline. For Blink, the $3\times 3$ variant offers a small gain in cognition but still reduces perception and overall MME performance, while the $4\times 4$ setup leads to drops across all metrics. This may be because the current thresholds are derived using a simple proportional scaling rule, and more fine-grained tuning could further optimize performance for different patch granularities. Nevertheless, across all configurations, Blink equipped with the trained TokenSR module consistently outperforms Blink-interp, confirming that the learned amplifier is more effective than naive interpolation under all patch partitions.
@@ Appendix F Ablation Study on Patch Numbers
Appendix G Ablation Study on Interpolation
Tab. 9 presents the performance on the MME benchmark with and without the interpolation step. In our experiments, the w/o interp condition refers to bypassing this initial upsampling. Specifically, after selecting the saliency tokens, they are directly extracted and inserted between the original visual and text tokens without resizing. For Blink-interp, w/o interp corresponds to inserting the saliency tokens directly into the dynamic inference pipeline without any training. For Blink, w/o interp means feeding the tokens into the trained TokenSR module without prior interpolation. From the experimental results, skipping interpolation for Blink-interp slightly reduces the perception score by 3.50, indicating that aligning the length of the expanded saliency sequence with the original image helps preserve spatial information. For Blink, removing interpolation leads to drops of 4.66 in perception, 6.79 in cognition, and 11.45 overall, further demonstrating that the TokenSR module benefits from receiving saliency tokens of the same length as the original image tokens, which allows the amplifier to more effectively refine multimodal representations. Overall, these results highlight that interpolation is a crucial step for fully leveraging spatial context and ensuring stable and accurate performance.
Appendix H Limitations and Future Work